# On measures derived from orbital integrals

### v2

@@ 3. The projection of an orbital measure to the Cartan algebra
For $a\in\operatorname{\mathfrak{t}}$ let $\nu_{a}$ be the orbital measure on $\operatorname{\mathcal{O}}_{a}$ and define the measure $\mu_{a}$ on the Cartan algebra as $\mu_{a}=\operatorname{pr}_{*}(\nu_{a})$ where $\operatorname{pr}:\operatorname{\mathfrak{g}}\to\operatorname{\mathfrak{t}}$ is the orthogonal projection. The measure $\mu_{a}$ was characterized in [Z16, O13, F15], it can also be characterized as a Duistermaat-Heckman measure in the context of Hamiltonian torus actions, see Section $5$ in the book [GLS96] and the references therein. We start with an equation satisfied by the measure $\mu_{a}$.
@@
(2) $\frac{\left[\Delta,\Delta\right]}{\Delta(a)\Delta(x)}\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)e^{\langle w(a),x\rangle}=\int_{\operatorname{\mathfrak{t}}}e^{\langle c,x\rangle}d\mu_{a}(c).$
@@
$\int_{G}e^{\langle\operatorname{Ad}_{g}a,x\rangle}dg=\int_{G}e^{\langle\operatorname{pr}(\operatorname{Ad}_{g}a),x\rangle}dg=\int_{\operatorname{\mathfrak{t}}}e^{\langle c,x\rangle}d\mu_{a}(c).$
@@
$\operatorname{\mathcal{F}}_{a}(x)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)\Delta(x)}\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)e^{\langle w(a),x\rangle}=\int_{\operatorname{\mathfrak{t}}}e^{\langle c,x\rangle}d\mu_{a}(c).$
@@
(3) $\int_{\operatorname{\mathfrak{t}}}f(c)d\mu_{a}(c)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)}\operatorname{ev}_{a}\left(I_{\Delta}P_{\operatorname{sym}}f\right),$
@@
We write $\eta=\left[\Delta,\Delta\right]/\Delta(a)$ and $l=|\Phi^{+}|$. If we take the part of degree not greater than $k\in\operatorname{\mathbb{N}}_{0}$ in $x$ of equation (2) of Proposition 3.1 then we have
$\eta\frac{r_{a}^{k+l}(x)}{\Delta(x)}=\int_{\operatorname{\mathfrak{t}}}q_{c}^{k}(x)d\mu_{a}(c).$
@@
where we used Propositions 2.9 in the second equality, Lemma 2.3 in the third and Proposition 2.4 in the final equality. On the other hand, in the LHS we get
$\left[\int_{\operatorname{\mathfrak{t}}}q_{c}^{k}d\mu_{a}(c),f_{\operatorname{sym}}\right]$ $=\int_{\operatorname{\mathfrak{t}}}\left[q_{c}^{k},f_{\operatorname{sym}}\right]d\mu_{a}(c)$
$=\int_{\operatorname{\mathfrak{t}}}\operatorname{ev}_{c}(f_{\operatorname{sym}})d\mu_{a}(c)$
$=\int_{\operatorname{\mathfrak{t}}}f_{\operatorname{sym}}(c)d\mu_{a}(c)$
$=\int_{\operatorname{\mathfrak{t}}}f(c)d\mu_{a}(c)$
@@
$\int_{\operatorname{\mathfrak{t}}}1d\mu(c)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)}\operatorname{ev}_{a}\left(I_{\Delta}\mathbf{1}\right)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)}\operatorname{ev}_{a}\left(\frac{\Delta}{\left[\Delta,\Delta\right]}\right)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)}\frac{\Delta(a)}{\left[\Delta,\Delta\right]}=1.$
@@
Consider the case in which $G=\operatorname{SU}(2)$, and $\alpha\in\operatorname{\mathfrak{t}}\subseteq\operatorname{\mathfrak{su}}(2)$ is given by $\alpha=e_{2}-e_{1}$ where $\{e_{1},e_{2}\}$ is the standard basis. Note that $\alpha^{\circ}=\frac{1}{2}\alpha$ and that $I_{\Delta}=I_{\alpha}$. To characterize the measure $\mu_{a}$ we compute
$\int_{\operatorname{\mathfrak{t}}}f(c)d\mu_{a}(c)$ $=\eta\operatorname{ev}_{a}\left(I_{\Delta}P_{\operatorname{sym}}f\right)$
$=\eta\operatorname{ev}_{a}\left(\int^{\langle x,\alpha^{\circ}\rangle}_{0}P_{\operatorname{sym}}f(x-\alpha t)dt\right)$
$=\eta\int^{\left\langle a,\frac{1}{2}\alpha\right\rangle}_{0}P_{\operatorname{sym}}f(a-\alpha t)dt$
$=\frac{2}{a_{2}-a_{1}}\int^{\frac{1}{2}(a_{2}-a_{1})}_{0}\frac{1}{2}\left(f(a-\alpha t)+f(-a+\alpha t)\right)dt.$
@@ 4. The radial part of the convolution of orbital measures
$=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(c)}\left(\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)e^{\langle w(c),x\rangle}\right)d\nu_{a,b}(c).$
@@
$\operatorname{\mathcal{F}}_{a}(z)\operatorname{\mathcal{F}}_{b}(z)=\int_{\operatorname{\mathfrak{t}}}\operatorname{\mathcal{F}}_{c}(z)d\nu_{a,b}(c).$
@@
(5) $\int_{\operatorname{\mathfrak{t}}}\frac{f_{\operatorname{alt}}(c)}{\Delta(c)}d\nu_{a,b}(c)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)\Delta(b)}\operatorname{ev}_{b}\left(I_{\Delta}\left(\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)T_{w(a)}f_{\operatorname{alt}}\right)\right).$
@@
$\varphi.F^{k}M_{r_{a}^{k}}\frac{r_{b}^{k+l}(x)}{\Delta(x)}=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(c)}r_{c}^{k}(x)d\nu_{a,b}(c).$
@@
where we used the definition $r_{a}^{k}=P_{\operatorname{alt}}q_{a}^{k}$ in the first equality, Proposition 2.5 in the third, Proposition 2.9 in the fifth, Lemma 2.3 in the sixth and Proposition 2.2 in the final equality. On the other hand, on the RHS we get
$\left[\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(c)}r_{c}^{k}(x)d\nu_{a,b}(c),f_{\operatorname{alt}}\right]$ $=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(c)}\left[r_{c}^{k}(x),f_{\operatorname{alt}}\right]d\nu_{a,b}(c)$
$=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(c)}\operatorname{ev}_{c}(f_{\operatorname{alt}})d\nu_{a,b}(c)$
$=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(c)}f_{\operatorname{alt}}(c)d\nu_{a,b}(c),$
@@
$\int_{\operatorname{\mathfrak{t}}}g(c)d\nu_{a,b}(c)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)\Delta(b)}\operatorname{ev}_{b}\left(I_{\Delta}\left(\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)T_{w(a)}(\Delta.P_{\operatorname{sym}}g)f\right)\right)$
@@
$\int_{\operatorname{\mathfrak{t}}}(P_{\operatorname{sym}}h)(c)d\nu_{a,b}(c)=\int_{\operatorname{\mathfrak{t}}}h(c)d\nu_{a,b}(c)$
@@
Consider the case in which $G=\operatorname{SU}(2)$, and $\alpha\in\operatorname{\mathfrak{t}}\subseteq\operatorname{\mathfrak{su}}(2)$ is given by $\alpha=e_{2}-e_{1}$ where $\{e_{1},e_{2}\}$ is the standard basis. Note that $\alpha^{\circ}=\frac{1}{2}\alpha$. For $f_{\operatorname{alt}}\in\operatorname{\mathcal{P}}_{\operatorname{alt}}(\operatorname{\mathfrak{t}})$
$\int_{\operatorname{\mathfrak{t}}}f_{\operatorname{alt}}(c)\frac{1}{\Delta(c)}d\nu_{a,b}(c)$ $=\varphi.\operatorname{ev}_{b}\left(I_{\Delta}\left(\frac{1}{2}\left(T_{a}f_{\operatorname{alt}}-T_{-a}f_{\operatorname{alt}}\right)\right)\right)$
$=\varphi.\operatorname{ev}_{b}\left(\int^{\langle x,\alpha^{\circ}\rangle}_{0}\frac{1}{2}\left((f_{\operatorname{alt}}(x+a-\alpha t)-f_{\operatorname{alt}}(x-a-\alpha t)\right)dt\right)$
$=\varphi.\int^{\frac{1}{2}(b_{2}-b_{1})}_{0}\frac{1}{2}\left(f_{\operatorname{alt}}(b+a-\alpha t)-f_{\operatorname{alt}}(b-a-\alpha t)\right)dt$
where $\varphi=\frac{2}{(a_{2}-a_{1})(b_{2}-b_{1})}$. If we make the identification $\operatorname{\mathfrak{t}}\simeq\operatorname{\mathbb{R}}$ given by $(x,-x)\to x$ then one can verify that the density function $\phi$ of the measure $\nu_{a,b}$ is given when $0<b<a$ by
$\phi(c)=\begin{cases}\frac{c}{4ab}&\text{if $a-b\leq c\leq a+b$}\\ \frac{-c}{4ab}&\text{if $-a-b\leq c\leq-a+b$}\\ 0&\text{otherwise.}\end{cases}$

### v3

@@ 3. The projection of an orbital measure to the Cartan algebra
For $a\in\operatorname{\mathfrak{t}}$ let $\nu_{a}$ be the orbital measure on $\operatorname{\mathcal{O}}_{a}$ and define the measure $\mu_{a}$ on the Cartan algebra as $\mu_{a}=\operatorname{pr}_{*}(\nu_{a})$ where $\operatorname{pr}:\operatorname{\mathfrak{g}}\to\operatorname{\mathfrak{t}}$ is the orthogonal projection. The measure $\mu_{a}$ was characterized in [Z16, O13, F15], it can also be characterized as a Duistermaat-Heckman measure in the context of Hamiltonian torus actions, see Section $5$ in the book [GLS96] and the references therein. In this section we give an alternative characterization and provide a formula for the moments in the case of type A root systems. We start with an equation satisfied by the measure $\mu_{a}$.
@@
(2) $\frac{\left[\Delta,\Delta\right]}{\Delta(a)\Delta(x)}\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)e^{\langle w(a),x\rangle}=\int_{\operatorname{\mathfrak{t}}}e^{\langle c,x\rangle}d\mu_{a}(x).$
@@
$\int_{G}e^{\langle\operatorname{Ad}_{g}a,x\rangle}dg=\int_{G}e^{\langle\operatorname{pr}(\operatorname{Ad}_{g}a),x\rangle}dg=\int_{\operatorname{\mathfrak{t}}}e^{\langle c,x\rangle}d\mu_{a}(x).$
@@
$\operatorname{\mathcal{F}}_{a}(x)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)\Delta(x)}\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)e^{\langle w(a),x\rangle}=\int_{\operatorname{\mathfrak{t}}}e^{\langle c,x\rangle}d\mu_{a}(x).$
@@
(3) $\int_{\operatorname{\mathfrak{t}}}f(x)d\mu_{a}(x)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)}\operatorname{ev}_{a}\left(I_{\Delta}P_{\operatorname{sym}}f\right),$
@@
We write $\eta=\left[\Delta,\Delta\right]/\Delta(a)$ and $l=|\Phi^{+}|$. If we take the part of degree not greater than $k\in\operatorname{\mathbb{N}}_{0}$ in the equation of Proposition 3.1 then we have
$\eta.\frac{r_{a}^{k+l}(x)}{\Delta(x)}=\int_{\operatorname{\mathfrak{t}}}q_{c}^{k}(x)d\mu_{a}(x).$
@@
where we used Definition 2.7 in the second equality, Lemma 2.3 in the third and Proposition 2.4 in the final equality. On the other hand, in the LHS we get
$\left[\int_{\operatorname{\mathfrak{t}}}q_{c}^{k}d\mu_{a}(x),f_{\operatorname{sym}}\right]$ $=\int_{\operatorname{\mathfrak{t}}}\left[q_{c}^{k},f_{\operatorname{sym}}\right]d\mu_{a}(x)$
$=\int_{\operatorname{\mathfrak{t}}}\operatorname{ev}_{c}(f_{\operatorname{sym}})d\mu_{a}(x)$
$=\int_{\operatorname{\mathfrak{t}}}f_{\operatorname{sym}}(x)d\mu_{a}(x)$
$=\int_{\operatorname{\mathfrak{t}}}f(x)d\mu_{a}(x)$
@@
Corollary 3.3.
We can rewrite (3) as
$\left(I_{\Delta}f\right)(a)=\frac{\Delta(a)}{\left[\Delta,\Delta\right]}\int_{\operatorname{\mathfrak{t}}}f(x)d\mu_{a}(x)$
for every $f\in\operatorname{\mathcal{P}}_{\operatorname{sym}}(\operatorname{\mathfrak{t}}).$ If we change the symbols $a$ and $x$ to $xx$ and $y$ we obtain
$\left(\Delta(\partial)^{-1}f\right)(x)=\left(I_{\Delta}f\right)(x)=\frac{\Delta(x)}{\left[\Delta,\Delta\right]}\int_{\operatorname{\mathfrak{t}}}f(y)d\mu_{x}(y).$
This provides an alternative characterization of $I_{\Delta}=\Delta(\partial)^{-1}$.
Remark 3.4.
In Proposition 2.10 we showed that
$I_{\Delta}(q_{b}^{k})=\frac{1}{\Delta(b)}r_{b}^{k+l}.$
Therefore if take $f=q_{b}^{k}$ in Theorem 3.2 we get
$\int_{\operatorname{\mathfrak{t}}}q_{a}^{k}(x)d\mu_{a}(x)$ $=\frac{\left[\Delta,\Delta\right]}{\Delta(a)}\operatorname{ev}_{a}\left(I_{\Delta}P_{\operatorname{sym}}q_{a}^{k}\right)$
$=\frac{\left[\Delta,\Delta\right]}{\Delta(a)\Delta(b)}r_{b}^{k+l}(a)$
which was the main assumption of the theorem.
@@
$\int_{\operatorname{\mathfrak{t}}}1d\mu(x)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)}\operatorname{ev}_{a}\left(I_{\Delta}\mathbf{1}\right)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)}\operatorname{ev}_{a}\left(\frac{\Delta}{\left[\Delta,\Delta\right]}\right)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)}\frac{\Delta(a)}{\left[\Delta,\Delta\right]}=1.$
@@
Consider the case in which $G=\operatorname{SU}(2)$. Using Proposition 2.8 and the identification $\operatorname{\mathfrak{t}}\simeq\operatorname{\mathbb{R}}$ given by $(x,-x)\mapsto x$ we have
$\int_{\operatorname{\mathfrak{t}}}f(x)d\mu_{a}(x)$ $=\eta.\operatorname{ev}_{a}\left(I_{\Delta}P_{\operatorname{sym}}f\right)$
$=\frac{1}{a}\int^{a}_{0}\frac{1}{2}\left(f(a-t)+f(-a+t)\right)dt.$
@@
We give a formula for the moments of the measure $\mu_{a}$ in the case of type A root systems.
Theorem 3.7.
In the case of the unitary group $\operatorname{U}(n)$ and for a partition $\eta\in\Pi$ the following equations for the moments of the monomial symmetric function $m^{\eta}(x)$ and the monomial $x^{\eta}$ hold
$\int_{\operatorname{\mathbb{R}}^{n}}m_{\eta}(x)d\mu_{b}(x)=\left[\Delta,\Delta\right]\sum_{\lambda\in\Pi}\frac{|S_{n}\cdot\eta|\eta!}{n!(\lambda+\delta)!}K_{\lambda\eta}s_{\lambda}(b),$
$\int_{\operatorname{\mathbb{R}}^{n}}x^{\eta}d\mu_{b}(x)=\left[\Delta,\Delta\right]\sum_{\lambda\in\Pi}\frac{\eta!}{n!(\lambda+\delta)!}K_{\lambda\eta}s_{\lambda}(b).$
Proof.
For a monomial symmetric polynomial $m_{\eta}$ Theorem 3.2 and Proposition 2.14 imply that
$\int_{\operatorname{\mathbb{R}}^{n}}m_{\eta}(x)d\mu_{b}(x)$ $=\frac{\left[\Delta,\Delta\right]}{\Delta(b)}\operatorname{ev}_{b}\left(I_{\Delta}m_{\eta}\right)$
$=\frac{\left[\Delta,\Delta\right]}{\Delta(b)}\operatorname{ev}_{b}\left(\sum_{\lambda}\frac{|S_{n}\cdot\eta|\eta!}{n!(\lambda+\delta)!}K_{\lambda\eta}a_{\lambda+\delta}\right)$
$=\frac{\left[\Delta,\Delta\right]}{a_{\delta}(b)}\left(\sum_{\lambda}\frac{|S_{n}\cdot\eta|\eta!}{n!(\lambda+\delta)!}K_{\lambda\eta}a_{\lambda+\delta}(b)\right)$
$=\left[\Delta,\Delta\right]\sum_{\lambda}\frac{|S_{n}\cdot\eta|\eta!}{n!(\lambda+\delta)!}K_{\lambda\eta}s_{\lambda}(b),$
therefore the first equation of the theorem holds. Also, it is easy to check that for a monomial symmetric polynomial $m^{\eta}$
$P_{\operatorname{sym}}(x^{\eta})=\frac{1}{|S_{n}\cdot\eta|}m^{\eta}(x).$
Hence, the second equation of the theorem follows. ∎
@@ 4. The radial part of the convolution of orbital measures
$=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(x)}\left(\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)e^{\langle w(x),x\rangle}\right)d\nu_{a,b}(x).$
@@
$\operatorname{\mathcal{F}}_{a}(z)\operatorname{\mathcal{F}}_{b}(z)=\int_{\operatorname{\mathfrak{t}}}\operatorname{\mathcal{F}}_{c}(z)d\nu_{a,b}(x).$
@@
(5) $\int_{\operatorname{\mathfrak{t}}}\frac{f_{\operatorname{alt}}(x)}{\Delta(x)}d\nu_{a,b}(x)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)\Delta(b)}\operatorname{ev}_{b}\left(I_{\Delta}\left(\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)T_{w(a)}f_{\operatorname{alt}}\right)\right).$
@@
$\varphi.F^{k}M_{r_{a}^{k}}\frac{r_{b}^{k+l}(x)}{\Delta(x)}=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(x)}r_{c}^{k}(x)d\nu_{a,b}(x).$
@@
where we used the definition $r_{a}^{k}=P_{\operatorname{alt}}q_{a}^{k}$ in the first equality, Proposition 2.5 in the third, Definition 2.7 in the fifth, Lemma 2.3 in the sixth and Proposition 2.2 in the final equality. On the other hand, on the RHS we get
$\left[\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(x)}r_{c}^{k}(x)d\nu_{a,b}(x),f_{\operatorname{alt}}\right]$ $=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(x)}\left[r_{c}^{k}(x),f_{\operatorname{alt}}\right]d\nu_{a,b}(x)$
$=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(x)}\operatorname{ev}_{c}(f_{\operatorname{alt}})d\nu_{a,b}(x)$
$=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(x)}f_{\operatorname{alt}}(x)d\nu_{a,b}(x),$
@@
$\int_{\operatorname{\mathfrak{t}}}g(x)d\nu_{a,b}(x)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)\Delta(b)}\operatorname{ev}_{b}\left(I_{\Delta}\left(\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)T_{w(a)}(\Delta.P_{\operatorname{sym}}g)f\right)\right)$
@@
$\int_{\operatorname{\mathfrak{t}}}(P_{\operatorname{sym}}h)(x)d\nu_{a,b}(x)=\int_{\operatorname{\mathfrak{t}}}h(x)d\nu_{a,b}(x)$
@@
Consider the case in which $G=\operatorname{SU}(2)$. For $f_{\operatorname{alt}}\in\operatorname{\mathcal{P}}_{\operatorname{alt}}(\operatorname{\mathfrak{t}})$, using the identification $\operatorname{\mathfrak{t}}\simeq\operatorname{\mathbb{R}}$ given by $(x,-x)\to x$ and Proposition 2.8 we have
$\int_{\operatorname{\mathfrak{t}}}f_{\operatorname{alt}}(x)\frac{1}{\Delta(x)}d\nu_{a,b}(x)$ $=\varphi.\operatorname{ev}_{b}\left(I_{\Delta}\left(\frac{1}{2}\left(T_{a}f_{\operatorname{alt}}-T_{-a}f_{\operatorname{alt}}\right)\right)\right)$
$=\frac{1}{2ab}\int^{b}_{0}\frac{1}{2}\left(f_{\operatorname{alt}}(b+a-t)-f_{\operatorname{alt}}(b-a-t)\right)dt$
One can verify that the density function $\phi$ of the measure $\nu_{a,b}$ is given when $0<b<a$ by
$\phi(x)=\begin{cases}\frac{x}{4ab}&\text{if $a-b\leq x\leq a+b$}\\ \frac{-x}{4ab}&\text{if $-a-b\leq x\leq-a+b$}\\ 0&\text{otherwise.}\end{cases}$